2. Calculate E for the titration of 48.00 mL of 0.110 M Ag when 50.00 mL of 0.140 M Cl is added. Silver wire is the indicato please answer the following question
Answer #1

2) According to the condition given in question moles of Agt = concentration x volume = 0.11M X 48,0 me = 5.28 mmol moles ofE = E-E- = [0.7993 – 0.0591 log+] -0.197 = 0.6023 +0.059) log [Agt ] we know that Ksp = [Ag+ ][cé ] [Aq+J = Ksp 118x10 ° 1 [c

.